OpenStudy (anonymous):

RECTANGLE PQRS HAS ITS COORDINATE AT P(-8,2), Q(-5,6), R(30) AND S(0,-4) A.)FIND ITS AREA B.)FIND ITS PERIMETER PLEASE I REALLY NEED YOUR HELP..

OpenStudy (lgbasallote):

First, graph it so you can visualize what it looks like. Soon, you find out that the length is line segmentQR or SQ. And the width is line segment PQ or RS. Get the length via distance formula. sqrt(x2 - x1)^2 + (y2 - y1)^2 length = sqrt of (3+5)^2 + (-6)^2 = sqrt 64 + 36 lenth = 10 units. width = sqrt (-5 +8)^2 +(6-2)^2 = sqrt (9 + 16) = sqrt (25) width = 5 Area = lw = (5)(10) = 50 sq units Perimeter = 2l + 2w = 2(5) + 2(10) = 10 + 20 = 30 units